HETAV PATEL

Grade 11,

A particle is projected with velocity U along x axis. The acceleration of the particle is proportional to the square of the distance from the origin a=kx², the distance at which the particle stops is A)√3u/2k B)(3u/2k)⅓ C)√2²/3k D)(3u²/2k)⅓

A particle is projected with velocity U along x axis. The acceleration of the particle is proportional to the square of the distance from the origin a=kx², the distance at which the particle stops is
A)√3u/2k
B)(3u/2k)⅓
C)√2²/3k
D)(3u²/2k)⅓

Grade:11

1 Answers

Arun
5 years ago
 
a = udu/dx = k*x^2 
udu = (k*x^2)dx
Integerating both sides,
 
k^2/2 = u*x^3/3
x = 3u/2k)^1/3
 
Hope it helps
 
Regards
Arun